OpenStudy (anonymous):

WILL GIVE MEDAL AND POTENTIALLY PAY!!! A quadratic equation is shown below: x2 − 8x + 13 = 0 Which of the following is the first correct step to write the above equation in the form (x − p)2 = q, where p and q are integers?

OpenStudy (anonymous):

This technique is called COMPLETING THE SQUARE If I write the equation in the form ax^2 + bx + c First, divide b in half (In this case, (-8)/2 = (-4)) This gives you - p (x -4))^2 Solve the above (x - b/2) (In this case, x^2 - 8x +16) You need to subtract 3 to get the real expression, so (x-4)^2 - 3 Now, to get it into the right format, just move the negative 3... (x-4)^2 = 3 Done! P = -4 Q = 3

OpenStudy (anonymous):

* P = 4

OpenStudy (anonymous):

these are my choices Which of the following is the first correct step to write the above equation in the form (x − p)2 = q, where p and q are integers? Subtract 5 from both sides of the equation Subtract 3 from both sides of the equation Add 5 to both sides of the equation Add 3 to both sides of the equation

OpenStudy (anonymous):

@livias.random

OpenStudy (anonymous):

Ugh. I did all that work for nothing? Whatever Anyways... add 3?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

are you sure its add 3 and not subtract 3? @livias.random